# Exploring the Impact of Nuclear Parton Distribution Functions on Open Heavy Flavor Azimuthal Angular Correlations in Pb-Pb Collisions at $\sqrt{s_{NN}}$ = 5.5 TeV

### v1

@@ IV.1 $\rm{D^{0}-\bar{D}^{0}}$ azimuthal angular correlation
    Table 1: Separation between two near-side peaks for different associated $p_{\rm{T}}$ ranges for trigger range 2 $<$ $p_{\rm{T}}^{\rm trig}$ $<$ 4 GeV/$c$.
  3. 3.
    FIG 3 shows a quantitative comparison between the width and yield of the NS peak. The lower panel of FIG 3 displays the widths of the two individual NS peaks ($\sigma_{1},\sigma_{2}$) and the combined width ($\sigma$). The individual widths of both peaks using free proton PDFs are consistent with the NS widths using nPDFs. However, the combined width is significantly higher than the yield obtained using the nPDFs. The combined width decreases with the inclusion of high $p_{\rm{T}}^{\rm assoc}$, which has a higher initial boost during production, resulting in a narrower opening angle.
  4. 4.
    Quantitatively, FIG 3 shows that for $p_{\rm{T}}^{\rm trig}$ Class I, the yields using proton PDFs and nPDFs are consistent. A small discrepancy in high $p_{\rm{T}}^{\rm assoc}$ can be attributed to uncertainties in PDFs and baseline estimation.
  5. 5.
    For $p_{\rm{T}}^{\rm trig}$ Class II and III, yields obtained using EPPS21 and nNNPDF3.0 are consistent, but the yield obtained using NNPDF2.3 is much lower. As $p_{\rm{T}}^{\rm assoc}$ grows, the discrepancy between free proton PDF and nPDF calculations rises.
  6. 6.
    For $p_{\rm{T}}^{\rm trig}$ Class I, the AS width computed using free proton PDFs is much smaller than the widths calculated using nPDFs widths. The difference between widths computed using proton PDF and nPDFs reduces with increasing $p_{\rm{T}}^{\rm assoc}$ and $p_{\rm{T}}^{\rm trig}$ ranges. Except at low $p_{\rm{T}}^{\rm assoc}$, the AS yield for proton PDFs and nPDFs is consistent.
  7. 7.
    The width of the AS peak also shows a decreasing trend for all ranges of $p_{\rm{T}}^{\rm assoc}$ and $p_{\rm{T}}^{\rm trig}$ in the case of nPDFs. This can be due to the increasing boost of associated particles with increasing $p_{\rm{T}}^{\rm assoc}$ and $p_{\rm{T}}^{\rm trig}$.
  8. 8.
    FIG 12 indicates that except for $p_{\rm{T}}^{\rm trig}$ Class I and $p_{\rm{T}}^{\rm assoc}$ between $\rm{2-4\,GeV/c}$, NS yields due to nPDFs are always bigger than free proton PDF (more than 30%). NS correlation widths are similar for $p_{\rm{T}}^{\rm trig}$ Class II and III across different PDFs, however for $p_{\rm{T}}^{\rm trig}$ Class I widths obtained using PDFs are narrower compared to free proton PDFs (more than 20%). On the other hand, FIG 13 shows that except for $p_{\rm{T}}^{\rm trig}$ Class I, nPDF NS yields are always smaller than free proton PDF (approximately 50%). AS correlation widths are compatible for $p_{\rm{T}}^{\rm trig}$ Class II and III among different PDFs, but for $p_{\rm{T}}^{\rm trig}$ Class I nPDF gives a broader correlation compared to free proton PDFs (more than 15%).
  9. 9.
@@
    The panel with $p_{\rm{T}}^{\rm trig}$ $\mathrm{>4\,GeV/c}$ and $\mathrm{4>}$ $p_{\rm{T}}^{\rm assoc}$ $\mathrm{>2\,GeV/c}$ is critical for measuring the degree of thermalization of the charm quark, because collective effects are dominant in this $p_{\rm{T}}$ range. In this range, medium effects can flatten the $\rm{D^{0}-\bar{D}^{0}}$ azimuthal angular correlation in actual Pb-Pb collisions. FIG 3 and 4 illustrate that the NS and AS peaks have identical widths within PDF uncertainties. nPDFs have no appreciable impact on the broadening of the azimuthal correlation in this $p_{\rm{T}}$ range for minimum bias Pb-Pb events. Central collisions are likely to provide more intriguing effects.
IV.2 $\rm{B-\bar{B}}$ meson azimuthal angular correlation
FIG 5 shows the differential azimuthal angular correlation of $\rm{B-\bar{B}}$ across all species of beauty mesons. The $p_{\rm{T}}^{\rm trig}$ ranges are $\mathrm{10-15}$,  $\mathrm{\geq\,15}$, and  $\mathrm{\geq\,20\,GeV/}c$, while the associated particles are picked in the $p_{\rm{T}}^{\rm assoc}$ ranges $\mathrm{5-10}$, $\mathrm{10-15}$, and $\mathrm{>5}$ GeV/$c$. FIG 7 and FIG 8 provide a quantitative comparison of yield and width of the NS and AS peaks of the azimuthal angular distribution, respectively. At low trigger and association $p_{\rm{T}}$ ranges, the away-side distribution (pair creation) dominates the near-side distribution (gluon splitting) due to its broader width. The choice of higher $p_{\rm{T}}$ ranges compared to $\rm{D^{0}-\bar{D}^{0}}$ correlation is made to study these features of azimuthal angular correlation arising due to production processes.
  1. 1.
    Qualitatively, the NS azimuthal angular distribution of $\rm{B-\bar{B}}$ is consistent with the nPDFs and free proton PDF. The discrepancy between proton PDF and nPDF is observed for the AS peak of correlation for $p_{\rm{T}}^{\rm trig}$ ranges between $\mathrm{10}<$$p_{\rm{T}}^{\rm trig}$ $<\mathrm{15}$ and $\mathrm{15}<$ $p_{\rm{T}}^{\rm trig}$. The near side peak of $\rm{B-\bar{B}}$ azimuthal correlation is much smaller than the away side peak, this is because the ratio of gluon splitting production cross-section to pair creation cross-section is smaller for beauty quark as compared to charm quark ratio Norrbin:2000zc .
    $p_{\rm{T}}^{\rm assoc}$ (GeV/$c$) Mean of near side peaks (rad)

### v2

@@ IV Results
Table 1: Angular separation between near-side (NS) peaks of $\rm{D^{0}-\bar{D}^{0}}$ correlation distribution for different associated $p_{\rm{T}}^{\rm\bar{D^{0}}}$ ranges and $p_{\rm{T}}^{\rm D^{0}}$ using NNPDF2.3 PDFs.
FIG. 2 further illustrates that, except for EPPS21 (Default MPI), all PDFs give a double-peak structure at NS of correlation distribution in the $p_{\rm{T}}^{\rm trig}$ range $2<$$p_{\rm{T}}^{\rm trig}$ $<4$ GeV/$c$. This structure arises from $\mathrm{c-\bar{c}}$ production via the gluon splitting process. In FIG. 3, a similar double-peak feature is observed for the $\rm{B-\bar{B}}$ correlation distribution across all $p_{\rm{T}}^{\rm assoc}$ classes. FIG. 10 (Appendix A) shows NS of $\rm{B-\bar{B}}$ correlation distribution for better visualization of double-peak.
Additionally, it is observed that the angle between the produced $\mathrm{c-\bar{c}}$ ($\mathrm{b-\bar{b}}$) pairs in the gluon splitting process decreases as $p_{\rm{T}}^{\rm assoc}$ increases within a given $p_{\rm{T}}^{\rm trig}$ range, as shown in TABLE 1 and TABLE 4 (Appendix A). This behavior can be explained by the fact that, at low $p_{\rm{T}}^{\rm assoc}$, the gluon splitting opening angle is wider, while at higher $p_{\rm{T}}^{\rm assoc}$, the angle narrows due to gluon boost.
$p_{\rm{T}}^{\rm trig}$ (GeV/$c$) Class I Class II Class III
$p_{\rm{T}}^{\rm assoc}$ (GeV/$c$) 2-4 2-6 $>$2 2-4 2-6 $>$2 2-4 2-6 $>$2
% change in near side yield with respect to NNPDF2.3
PDF Settings
EPPS21+NNPDF2.3 -15 -9 -6 -21 -21 -20 -10 -11 -14
EPPS21+CT18ANLO -55 -66 -34 -38 -29 -27 1 -2 -12
nNNPDF3.0+NNPDF4.0 -82 -73 -54 -4 -16 -15 3 -4 -4
EPPS21 -90 -77 -74 -17 -18 -18 1 -2.0 -3
EPPS21 (Default MPI) -21 29 41 62 35 27 40 26 17
% change in near side width with respect to NNPDF2.3
EPPS21+NNPDF2.3 -1 0 0 -10 -12 -11 -10 -9 -8
EPPS21+CT18ANLO 18 4 4 -19 -15 -13 -5 0 -4
nNNPDF3.0+NNPDF4.0 -13 -28 -20 -10 -12 -20 -3 -9 -6
EPPS21 -49 -34 -34 -16 -15 -14 -6 -7 -6
EPPS21 (Default MPI) -43 -28 -23 -10 -12 -11 -1 -4 -6
Table 2: Impact of different PDF settings on the near-side (AS) yield and width of $\rm{D^{0}-\bar{D}^{0}}$ azimuthal angular correlations in Pb–Pb collisions at $\sqrt{s_{NN}}=5.5$ TeV. Values indicate the percentage deviation from the NNPDF2.3 baseline, presented for various $p_{\rm{T}}^{\rm trig}$ classes and $p_{\rm{T}}^{\rm assoc}$ ranges.
$p_{\rm{T}}^{\rm trig}$ (GeV/$c$) Class I Class II Class III
$p_{\rm{T}}^{\rm assoc}$ (GeV/$c$) 2-4 2-6 $>$2 2-4 2-6 $>$2 2-4 2-6 $>$2
% change in away side yield with respect to NNPDF2.3
PDF Settings
EPPS21+NNPDF2.3 -14 -4 -4 -3 -4 -4 -1 -2 -3
EPPS21+CT18ANLO -40 -34 -22 -10 -13 -15 -4 -8 -11
nNNPDF3.0+NNPDF4.0 -19 -7 4 3 -6 -7 1 -5 -7
EPPS21 92 17 1 -1 -10 -12 -4 -9 -11
EPPS21 (Default MPI) 220 100 72 11 1 -1 -2 -8 -8
% change in away side width with respect to NNPDF2.3
EPPS21+NNPDF2.3 -7 -2 -1 -2 -2 -3 -1 -2 -3
EPPS21+CT18ANLO 34 12 16 4 5 4 4 4 2
nNNPDF3.0+NNPDF4.0 36 26 21 8 6 6 3 3 3
EPPS21 41 34 25 7 6 6 2 3 3
EPPS21 (Default MPI) 21 11 11 0 -1 -1 -1 -1 -1
Table 3: Impact of different PDF settings on the away-side (AS) yield and width of $\rm{D^{0}-\bar{D}^{0}}$ azimuthal angular correlations in Pb–Pb collisions at $\sqrt{s_{NN}}=5.5$ TeV. Values indicate the percentage deviation from the NNPDF2.3 baseline, presented for various $p_{\rm{T}}^{\rm trig}$ classes and $p_{\rm{T}}^{\rm assoc}$ ranges.
FIG. 4 and FIG. 6 present the yield and width of the NS distribution for $\rm{D^{0}}$ and B-meson correlations, respectively, while FIG. 5 and FIG. 7 depict the same for the AS distribution. For $\rm{D^{0}}$ mesons, the NS yield increases with both $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ across all PDF settings. FIG. 4 shows that, the NS width decreases as a function of $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ except in the case of EPPS21. For $\rm{B-\bar{B}}$ correlations (FIG. 6), the NS width exhibits minimal $p_{\rm{T}}^{\rm assoc}$ dependence in $p_{\rm{T}}^{\rm trig}$ Classes II and III, whereas in $p_{\rm{T}}^{\rm trig}$ Class I, it increases with $p_{\rm{T}}^{\rm assoc}$. For $p_{\rm{T}}^{\rm assoc}$ $>5$ GeV/$c$, the NS width decreases with $p_{\rm{T}}^{\rm trig}$ due to the greater boost of high $p_{\rm{T}}$ trigger and associated particles. However, for $5<$ $p_{\rm{T}}^{\rm assoc}$ $<10$ GeV/$c$, the NS width increases with $p_{\rm{T}}^{\rm trig}$. In this $p_{\rm{T}}^{\rm assoc}$ range, associated particles undergo significant fragmentation and scattering compared to the trigger particle, reducing the correlation strength and resulting in a broader angular distribution. A similar effect is observed for the AS width (FIG. 7), where $p_{\rm{T}}^{\rm trig}$ Class III exhibits a broader distribution than Class II. This effect is more pronounced in the AS than in the NS due to the smaller $\Delta p$ of the heavy quarks produced via the pair creation process compared to those from gluon splitting Norrbin:2000zc . However, with the inclusion of higher-$p_{\rm{T}}$ associated particles, the expected trend of decreasing width with $p_{\rm{T}}^{\rm trig}$ is restored.
FIG. 5 shows that the AS yield increases with $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ for $\rm{D^{0}}$ mesons. However, for B-mesons (FIG. 7), the AS yield rises in $p_{\rm{T}}^{\rm trig}$ classes I and II but decreases in Class III. The difference in AS yield between classes II and III diminishes as higher-$p_{\rm{T}}$ associated $\bar{B}$ mesons are included. For $p_{\rm{T}}^{\rm assoc}$ $\rm{>5}$ GeV/$c$, the AS yield in $p_{\rm{T}}^{\rm trig}$ Class III is higher than that of Class II, restoring the usual $p_{\rm{T}}^{\rm assoc}$ dependence. These observations can be attributed to the availability of associated particles for $p_{\rm{T}}^{\rm trig}$ Class III compared to Class II. Associated particles predominantly exhibit $\Delta p\approx 3~{}\mathrm{GeV}/c$ for pair creation and $\Delta p\approx 5~{}\mathrm{GeV}/c$ for gluon splitting Norrbin:2000zc . As a result, $p_{\rm{T}}^{\rm trig}$ Class III has fewer associated particles than $p_{\rm{T}}^{\rm trig}$ Class II in $p_{\rm{T}}^{\rm assoc}$ range 5-10 GeV/$c$. In contrast, the NS distribution does not follow this trend due to the larger $\Delta p$ range for the gluon splitting process. This behavior remains consistent across all PDFs and nPDFs under consideration.
As discussed earlier, the choice of nPDFs for all processes significantly reduces the multiplicity distribution due to a lower number of MPIs compared to the default PDF. EPPS21 with the number of MPIs same as default PYTHIA8 settings, exhibits a double-peak structure in the $p_{\rm{T}}^{\rm trig}$ range $2<$$p_{\rm{T}}^{\rm trig}$ $<4$ GeV/$c$, which is absent when using EPPS21 with default MPI settings. This suggests that the double-peak structure arises from hard MPIs enhancing $\rm{D^{0}}$ production at low $p_{\rm{T}}$, making the double peak more pronounced.
A comparison between EPPS21+NNPDF2.3, EPPS21+CT18ANLO, and nNNPDF3+NNPDF4 where former PDFs are used only for hard scatterings and later PDFs for ISR, FSR, and MPIs Bierlich:2022pfr , shows that proton PDF choice for MPI, ISR, and FSR significantly impacts correlation shape, yield, and width. The NS peak shape remains similar across PDFs, but correlation distribution is heavily suppressed for CT18ANLO and NNPDF4 in $p_{\rm{T}}^{\rm trig}$ Class I compared to NNPDF2.3. Quantitatively, TABLE 2 show that the NS yield decreases by 30–65% (CT18ANLO) and 50–80% (NNPDF4) depending on $p_{\rm{T}}^{\rm assoc}$ in $p_{\rm{T}}^{\rm trig}$ Class I. For $p_{\rm{T}}^{\rm trig}$ Class II, NS yield decreases by $\approx$ 25-40% for CT18ANLO, while for NNPDF4 yields do not change more than 20%. The CT18ANLO does not significantly alter the NS width of the correlation, but NNPDF4 exhibits a 20–30% narrower correlation for the lowest $p_{\rm{T}}^{\rm trig}$ range. The AS peak remains consistently smaller than the default PDF for CT18ANLO and NNPDF4 across all $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ ranges (FIG 2). For the $p_{\rm{T}}^{\rm trig}$ Class I, yield decreases up to 40% for CT18ANLO, while deviation remains below 20% for NNPDF4. The AS width increases by up to 35% for NNPDF4 and up to 30% for CT18ANLO as reported in TABLE 3.
A comparison between NNPDF2.3 and EPPS21 reveals that NS correlations are significantly suppressed, whereas AS correlations are larger for EPPS21 compared to NNPDF2.3 (FIG. 2). As reported in TABLE 2, the NS yield is 70-90% less than the default PDFs for $p_{\rm{T}}^{\rm trig}$ Class I. The AS yield is twice that of the default PDFs for $\rm{2-4\,GeV/c}$ but becomes comparable when including high $p_{\rm{T}}$ associated particles. For other $p_{\rm{T}}^{\rm trig}$ ranges, the difference remains below 10%. As per the TABLE 3, NS widths decrease, whereas AS widths increase by 20–50% for EPPS21 in the lowest $p_{\rm{T}}^{\rm trig}$ range, depending on $p_{\rm{T}}^{\rm assoc}$. In $p_{\rm{T}}^{\rm trig}$ Class II, NS widths deviate by more than 10% at the lowest $p_{\rm{T}}^{\rm assoc}$, whereas AS widths remain largely unchanged. These results highlight significant modifications due to the nuclear environment, particularly at low $p_{\rm{T}}^{\rm trig}$.
A comparison between NNPDF2.3 and EPPS21+NNPDF2.3 examines the impact of nPDFs on hard scattering in the correlation distribution. The NS and AS distributions remain similar across all $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ ranges, indicating that in PYTHIA8, the choice of nPDFs for hard scattering does not affect the azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$. This suggests that the azimuthal angular distribution of $\rm{D^{0}-\bar{D}^{0}}$ is primarily governed by the PDFs used for ISR, FSR, and MPIs.
@@
The panel with $p_{\rm{T}}^{\rm trig}$ $>4$ GeV/$c$ and $2<$ $p_{\rm{T}}^{\rm assoc}$ $<4$ GeV/$c$ is critical for studying charm quark thermalization, as collective effects dominate in this range. In Pb–Pb collisions, medium effects can flatten the $\rm{D^{0}-\bar{D}^{0}}$ azimuthal angular correlation. TABLE 2 and TABLE 3 show that NS and AS peaks have identical widths within PDF uncertainties (less than 15%), indicating that nPDFs have no significant impact on azimuthal broadening in this $p_{\rm{T}}$ range for minimum bias Pb–Pb events.
In contrast to $\rm{D^{0}-\bar{D}^{0}}$ correlations, $\rm{B-\bar{B}}$ correlations do not show noticeable sensitivity to PDF variations. This can be due to higher $p_{\rm{T}}$ ranges of trigger and associated particles. These higher $p_{\rm{T}}$ values correspond to larger Bjorken-x, where PDFs are more consistent across parameterizations, reducing their impact on $\rm{B-\bar{B}}$ correlations.
Figure 10: Comparison of NS azimuthal angular correlation of $\rm{B-\bar{B}}$ meson between free proton PDF (NNPDF2.3) and nuclear PDF (EPPS21,nNNPDF3) with PYTHIA8+Angantyr in Pb-Pb collisions at $\rm{\sqrt{s}=5.5\,TeV}$ is shown for different $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ ranges. The NS peak is plotted for $\rm{|\Delta\phi|}$ = 1.
V Summary
In this study, we investigate the impact of state-of-the-art PDFs on the azimuthal angular correlation of $\rm{D^{0}-\bar{D}^{0}}$, and $\rm{B-\bar{B}}$. We simulate Pb-Pb collisions at $\rm{\sqrt{s}=5.5\,TeV}$ using the PYTHIA8 event generator and construct the azimuthal angular correlations of the open heavy flavor mesons by correlating the heavy quark and anti-quark pairs from the same hard scattering. By delving into this analysis, we seek to uncover potential nuclear influences on thermalization and the production of heavy quarks, both at low and high $p_{\rm{T}}$. The key findings of our investigation are outlined as follows:
  • A comparison between NNPDF2.3 and multiplicity-corrected EPPS21 implies that nuclear PDFs significantly modify the azimuthal angular correlation of open heavy-flavor mesons compared to free proton PDFs. These modifications influence both the yield and width of NS and AS correlations across all $p_{\rm{T}}^{\rm assoc}$ and $p_{\rm{T}}^{\rm trig}$ ranges by altering ISR, FSR, and MPIs.
  • The nuclear PDF effects vary with both trigger and associated $p_{\rm{T}}$. These effects are most pronounced at low $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$, where the sampled Bjorken-$x$ values are small, and nuclear modifications such as shadowing play a significant role. At high $p_{\rm{T}}$, the nuclear modifications become less significant as PDFs converge at large Bjorken-$x$.
  • A double-peak structure is observed at the NS in $\rm{D^{0}-\bar{D}^{0}}$ and $\rm{B-\bar{B}}$ azimuthal correlations, providing a clear signature of partonic heavy quark production via the gluon splitting process. The separation between the peaks decreases with increasing $p_{\rm{T}}^{\rm assoc}$ due to a stronger initial boost.
  • A comparison between EPPS21 and EPPS21 (Default MPI) reveals that increased hard multi-partonic interactions (MPIs) in nPDFs enhance $\rm{D^{0}}$ production at low $p_{\rm{T}}$, making the double-peak structure more prominent. This highlights the role of MPI in shaping the azimuthal angular distribution in high-energy nuclear collisions.
  • The observed modifications in azimuthal angular correlations arise from the choice of nuclear PDFs rather than collective effects. The $p_{\rm{T}}^{\rm trig}$ $>4$ GeV/c and $2<$ $p_{\rm{T}}^{\rm assoc}$ $<4$ GeV/$c$ range is crucial for studying charm quark thermalization. Nuclear PDFs alone do not alter the width of the azimuthal correlation significantly in minimum bias Pb–Pb events for these $p_{\rm{T}}$ ranges. For lower $p_{\rm{T}}$ ranges, the impact of nPDFs needs to be taken into account.
  • This study introduces a new fitting function for the azimuthal angular correlation of $\rm{B-\bar{B}}$ mesons, formulated as a convolution of Gaussian and triangular functions. Unlike previous models based on generalized Gaussian or von Mises distributions, this function more accurately captures the evolution of the AS peak from Gaussian-like at low $p_{\rm{T}}$ to triangular-like at high $p_{\rm{T}}$. This improved parametrization provides a better description of beauty quark correlation dynamics and could be useful for future studies.
  • Unlike $\rm{D^{0}-\bar{D}^{0}}$ correlations, nuclear PDFs have a weaker effect on $\rm{B-\bar{B}}$ azimuthal correlations. This may be due to the higher mass of the beauty quark, which reduces sensitivity to initial-state nuclear modifications, or the higher $p_{\rm{T}}$ ranges used in $\rm{B-\bar{B}}$ studies. This suggests that nuclear PDF effects can be overlooked for beauty quark studies at high $p_{\rm{T}}$ but remain important for charm.
  • Differential study conducted in this article shows that kinematics of heavy flavor production can be studied in greater detail using azimuthal angular correlations.
A detailed study of the azimuthal angular correlation of open heavy flavors provides valuable insights into heavy flavor dynamics in both small and large systems. The findings of this study, based on minimum bias Pb–Pb events, can be extended to centrality-dependent analyses, which may reveal additional nuclear modification effects and offer deeper insights into the complex dynamics of heavy quark production and evolution.
Appendix A Near side features of $\rm{B-\bar{B}}$ correlation
To further investigate the NS structure of $\rm{B-\bar{B}}$ azimuthal correlations, we analyze the peak separation for different PDFs in Pb-Pb collisions at $\sqrt{s}=5.5$ TeV using the PYTHIA8+Angantyr framework. FIG. 10 presents the NS azimuthal angular correlation of $\rm{B-\bar{B}}$ mesons for various $p_{\rm{T}}^{\rm trig}$ and $p_{\rm{T}}^{\rm assoc}$ ranges, highlighting differences between free proton PDFs (NNPDF2.3) and nuclear PDFs (EPPS21, nNNPDF3). The NS peak is examined within $\rm{|\Delta\phi|}=1$. Table 4 quantifies the separation between the two NS peaks across different kinematic classes. The observed trend suggests that peak separation decreases as $p_{\rm{T}}^{\rm trig}$ increases, indicating a more collimated correlation due to the larger boost of associated and trigger particles at higher momenta.
$p_{\rm{T}}^{\rm assoc}$ (GeV/$c$) Mean of NS peaks (rad)